# Self-Normalized Resets for Plasticity in Continual Learning

### v1

@@ 1 Introduction
One formal definition of plasticity measures the ability of a network initialized at a specific set of parameters to fit a random target function using some pre-specified optimization procedure. In this sense, random parameter initializations (eg. Lyle et al. (2024)) are known to enjoy high plasticity. This has motivated two related classes of algorithms that attempt to mitigate plasticity loss. The first explicitly ‘resets’ neurons that are deemed to have low ‘utility’ (Dohare et al., 2023; Sokar et al., 2023). A reset re-initializes the neurons input weights and bias according to some suitable random initialization rule, and sets the output weights to zero; algorithms vary in how the utility of a neuron is defined and estimated from online data. A second class of algorithms perform this reset procedure implicitly via regularization (Ash & Adams, 2020; Kumar et al., 2023). These latter algorithms differ in their choice of what to regularize towards, with choices including the original network initialization; a new randomly drawn initialization; or even zero. The aforementioned approaches to mitigating plasticity loss attempt to adjust the training process; other research has studied the role of architectural and optimizer hyper-parameter choices. Across all of the approaches to mitigating plasticity loss described above, no single approach is yet to emerge as both robust to hyper-parameter choices, and simultaneously performant across benchmark problems.
Given some point process consider the task of distinguishing between the hypotheses that this point process has a positive rate (the null hypothesis), or a rate that is identically zero with a penalty for late rejection or acceptance. An optimal test here takes the following simple form: we reject the null hypothesis as soon as the time elapsed without an event exceeds some percentile of the inter-arrival time under the null hypothesis and otherwise accept immediately upon an event. Viewing the firing of a neuron as such a point process, we propose to reset a neuron based on a rejection of the hypothesis that the the neuron is firing at a positive rate. We use the histogram of past inter-firing times as a proxy of the inter-arrival time distribution under the null hypothesis. This exceedingly simple algorithm is specified by a single hyper-parameter: the rejection percentile threshold. We refer to this procedure as self-normalized resets (SNR) and argue this is a promising approach to mitigating plasticity loss:
@@
    We demonstrate superior performance on four benchmark problems classes studied in (Dohare et al., 2023; Kumar et al., 2023). Interestingly, there is no single closest competitor to SNR across these problems. Many competing approaches also show significant sensitivity to the choice of hyper-parameters; SNR does not. We introduce a new problem to elucidate similar plasticity loss phenomena in the context of language models, and show similar relative merits for SNR.
@@ 1.1 Related Literature
Regularization Methods. L2 regularization has been shown to reduce plasticity loss, but is insufficient in completely alleviating the phenomenon (Dohare et al., 2021; Lyle et al., 2023). While L2 regularization limits weight norm growth during continual learning, it can exacerbate weight rank collapse due to regularization towards the origin. One successful regularization technique is Shrink and Perturb (S&P) (Ash & Adams, 2020), which periodically scales the network’s weights by a shrinkage factor $p$ followed by adding random noise to each weight with scale $\sigma$. Another approach is to perform L2 regularization towards the initial weights referred to as L2 Init (Kumar et al., 2023). These methods can be viewed as variants of L2 regularization that regularize towards a random initialization and the original initialization, respectively. These methods limit the growth of weight norms while maintaining weight rank and neuron activity by regularizing towards a high-plasticity parameterization.
Architectural and Optimizer Modifications. Architectural modifications such as layer normalization (Ba et al., 2016) and the use of concatenated ReLU activations have been shown to improve plasticity to varying degrees across network architectures and problem settings (Lyle et al., 2023; Kumar et al., 2023). Additionally, tuning Adam hyperparameters to improve the rate at which second moment estimates are updated has been explored with some success in Lyle et al. (2023).
@@ 2 Algorithm
Plasticity loss describes the phenomenon where, for certain continual learning processes $\Theta_{t}$, such as those corresponding to SGD or Adam, average expected regret increases over time, even for benign choices of $\mu_{t}$111such as a sequence corresponding to multiple epochs on a random sub-sample of a dataset followed by multiple epochs on a second random sub-sample; or Example 2.1. To make these ideas concrete, it is worth considering an example of the above phenomenon reported first by Dohare et al. (2021).
@@
Consider a sequence of ‘tasks’ presented sequentially to SGD, wherein each task consists of 10000 images from the MNIST dataset with the pixels permuted. SGD trains over a single epoch on each task before the subsequent task is presented. Figure 1 measures average accuracy on each task; we see that average accuracy decreases over tasks. The figure also shows a potential correlate of this phenomenon: the number of ‘dead’ or inactive neurons in the network increases as training proceeds, diminishing the network’s effective capacity.
@@
To motivate our algorithm, SNR, consider applying the network $f(\cdot;\theta_{t}^{*})$ to a hypothetical sequence of training examples drawn i.i.d. from $\mu_{t}$. Let $Z^{\mu_{t}}_{s,i}$ indicate the sequence of activations of neuron $i$, and let let $A^{\mu_{t}}_{i}$ denote the random time between two consecutive activations over this hypothetical sequence of examples. Now turning to the actual sequence of training examples, let $a^{t}_{i}$ count the time since the last firing of neuron $i$ prior to time $t$. Our (idealized) proposal is then exceedingly simple: reset neuron $i$ at time $t$ iff $\mathbb{P}(A_{i}^{\mu_{t}}\geq a^{t}_{i})\leq\eta$ for some suitably small threshold $\eta$. We dub this algorithm Self-Normalized Resets and present it as Algorithm 1. The algorithm requires a single hyper parameter, $\eta$. Of course in practice, the distribution of $A_{i}^{\mu_{t}}$ is unknown to us, and so an implementable version of Algorithm 1 simply approximates this distribution with the histogram of inter-firing times of neuron $i$ prior to time $t$ [^2] As opposed to tacking the histogram itself, we simply track the mean inter-firing time, and assume $A_{i}^{\mu_{t}}$ is geometrically distributed with that mean. This requires tracking just one parameter per neuron.
@@ 2.1 Motivating SNR and Comparison to Other Reset Schemes
Here we motivate the SNR heuristic and compare it to other proposed reset schemes. Consider the following simple hypothesis test: we observe a discrete time process $Z_{s}\in\{0,1\}$ which under the null hypothesis $H_{0}$ is a Bernoulli process with mean $p>0$. The alternative hypothesis $H_{1}$ is that the mean of the process is identically zero. A hypothesis test must, at some stopping time $\tau$, either reject ($X_{\tau}=1$) or accept ($X_{\tau}=0$) the null; an optimal such test would choose to minimize the sum of type-1 and type-2 errors and a penalty for delays:
@@
Here $\lambda>0$ penalizes the delay in a decision. If $\lambda<p/2$, the optimal test takes a simple form: for some suitable threshold $\bar{T}$, reject the null hypothesis iff $Z_{s}=0$ for all times $s$ up to $\bar{T}$:
@@
Notice that if $\lambda\propto p$, the percentile threshold above is independent of $p$. Applying this setup to the setting where under the null, we observe the firing of neuron $i$ under i.i.d. training examples from $\mu_{t}$, imagine that $p=\mathbb{P}(Z^{\mu_{t}}_{s,i}=1)$. Further, we assume $\lambda=\alpha p$; a reasonable assumption which models a larger penalty for late detection of neurons that are highly active. It is then optimal to declare neuron $i$ ‘inactive’ if the length of time it has not fired exceed the $1-\alpha(1-\alpha)^{-1}$ percentile of the distribution of $A_{i}^{\mu_{t}}$. This is the underlying motivation for the SNR heuristic.
@@
This reveals an interesting comparison with SNR. The schemes above will re-initialize a neuron after inactivity over a period of time that is uniform across all neurons. In the context of the hypothesis testing setup above, this will result in sub-optimal error rates across neurons. On the other hand, SNR will reset a neuron after it is inactive for a period that is effectively normalized to the nominal firing rate of that neuron, while still only specifying a single hyperparameter for the network.
@@ 3.1 Experimental Setup
Permuted MNIST (PM) (Goodfellow et al., 2013; Dohare et al., 2021; Kumar et al., 2023): A subset of 10000 image-label pairs from the MNIST dataset are sampled for an experiment. A task consists of a random permutation applied to each of the 10000 images. The network is presented with 2400 tasks appearing in consecutive order. Each task consists of a single epoch and the network receives data in batches of size 16.
Random Label MNIST (RM) (Kumar et al., 2023; Lyle et al., 2023): A subset of 1200 images from the MNIST dataset are sampled for an experiment. An experiment consists of 100 tasks, where each tasks is a random assignment of labels, consisting of 10 classes, to the 1200 images. A network is trained for 400 epochs on each task with a batch size 16.
Random Label CIFAR (RC) (Kumar et al., 2023; Lyle et al., 2023): A subset of 128 images from the CIFAR-10 dataset are sampled for an experiment. An experiment consists of 50 tasks, where each tasks is a random assignment of labels, consisting of 10 classes, to the 128 images. An agent is trained for 400 epochs on each task with a batch size 16.
Continual Imagenet (CI) (Dohare et al., 2023; Kumar et al., 2023): An experiment consists of all 1000 classes of images from the ImageNet-32 dataset (Chrabaszcz et al., 2017) containing 600 images from each class. Each task is a binary classification problem between two of the 500 classes, selected at random. The experiment consists of 500 tasks and each class occurs in exactly one task. Each task consists of 1200 images, 600 from each class, and the network is trained for 10 epochs with a batch size of 100.
@@
This experimental setup, for all but Permuted Shakespeare, follows that of (Kumar et al., 2023), with the exceptions of Permuted MNIST which has its task count increased from 500 to 2400, Random Label MNIST which has its task count increased from 50 to 100, and Random Label CIFAR which has its dataset reduced from 1200 to 128 images. Lyle et al. (2023) consider variants of the Random Label MNIST and CIFAR problems by framing them as MDP environments for DQN agents. During training, the DQN agents are periodically paused to assess their plasticity by training them on separate, randomly generated regression tasks using the same image datasets.
@@ 3.1.1 Algorithms and Architetcures
Our baseline in all problems consist simply of using SGD or Adam as the optimizer with no further intervention. We then consider several interventions to mitigate plasticity loss. First, we consider algorithms that employ an explicit reset of neurons: these include SNR, Continual Backprop (CBP) (Dohare et al., 2021), and ReDO (Sokar et al., 2023). Among algorithms that attempt to use regularization, we consider vanilla L2 regularization, L2 Init (Kumar et al., 2023), and Shrink and Perturb (Ash & Adams, 2020). Finally, as a potential architectural modification we consider the use of Layer Normalization (Ba et al., 2016).
@@
MLP: For Permuted MNIST and Random Label MNIST we use an MLP identical to that in Kumar et al. (2023) which in turn is a slight modification to that in Dohare et al. (2023).
CNN: For Random Label CIFAR and Continual ImageNet we use a CNN architectures identical to that in Kumar et al. (2023) which in turn is a slight modification to that in Dohare et al. (2023).
@@ 3.2 Results and Discussion
Table 1: Average accuracy on the last 10% of tasks for each algorithm on the benchmark continual learning problems Permuted MNIST (PM), Random Label MNIST (RM), Random Label CIFAR (RC), and Continual ImageNet (CI). Each algorithm is trained for 5 seeds, with both Adam and SGD, with its best performing hyperparameters after an initial hyperparameter sweep over 5 seeds.
@@
L2 $0.2762\pm 0.0309$ $0.1560\pm 0.0236$ $0.3101\pm 0.0425$
SNR+L2 $\textbf{0.2177}\pm 0.0196$ $\textbf{0.1370}\pm 0.0169$ $\textbf{0.2551}\pm 0.0454$
No Intv. $2.7397\pm 0.0140$ $1.8164\pm 0.0295$ $3.0147\pm 0.0250$
L2 Init $1.5052\pm 0.0437$ $1.2931\pm 0.0486$ $1.5262\pm 0.0420$
SNR $2.6872\pm 0.0222$ $1.7338\pm 0.0295$ $3.0242\pm 0.0408$
Table 2: Average loss measured on the final epoch of each task (mean $\pm$ standard deviation over 9 seeds) for different continual learning algorithms on the Permuted Shakespeare problem. Results are reported for the best performing hyperparameter(s) of each algorithm, selected after an initial hyperparameter sweep over 5 seeds.
@@
While L2 regularization addresses weight blowup, neuron death remains present; see Figure 3. The average loss with L2 increases from 0.1560, over the first 50 tasks, to 0.3101, over the final 50 tasks. This prompts us to consider using SNR in addition to L2 regularization. This largely eliminates neuron death (see the right panel of Figure 3), while stabilizing weight norms and maintaining entropy of self-attention probabilities, providing the lowest loss (0.2551) over the final 50 tasks.
Figure 3: The number of neurons that have been inactive for the last 5 consecutive batches on the Permuted Shakespeare problem, averaged over time steps within each epoch.

### v2

@@ 1 Introduction
One formal definition of plasticity measures the ability of a network initialized at a specific set of parameters to fit a random target function using some pre-specified optimization procedure. In this sense, random parameter initializations (eg. Lyle et al. (2024)) are known to enjoy high plasticity. This has motivated two related classes of algorithms that attempt to mitigate plasticity loss. The first explicitly ‘resets’ neurons that are deemed to have low ‘utility’ (Dohare et al., 2023; Sokar et al., 2023). A reset re-initializes the neurons input weights and bias according to some suitable random initialization rule, and sets the output weights to zero; algorithms vary in how the utility of a neuron is defined and estimated from online data. A second class of algorithms perform this reset procedure implicitly via regularization (Ash & Adams, 2020; Kumar et al., 2023b). These latter algorithms differ in their choice of what to regularize towards, with choices including the original network initialization; a new randomly drawn initialization; or even zero. The aforementioned approaches to mitigating plasticity loss attempt to adjust the training process; other research has studied the role of architectural and optimizer hyperparameter choices. Across all of the approaches to mitigating plasticity loss described above, no single approach is yet to emerge as both robust to hyperparameter choices, and simultaneously performant across benchmark problems.
Given some point process consider the task of distinguishing between the hypotheses that this point process has a positive rate (the null hypothesis), or a rate that is identically zero with a penalty for late rejection or acceptance. An optimal test here takes the following simple form: we reject the null hypothesis as soon as the time elapsed without an event exceeds some percentile of the inter-arrival time under the null hypothesis and otherwise accept immediately upon an event. Viewing the firing of a neuron as such a point process, we propose to reset a neuron based on a rejection of the hypothesis that the the neuron is firing at a positive rate. We use the histogram of past inter-firing times as a proxy of the inter-arrival time distribution under the null hypothesis. This exceedingly simple algorithm is specified by a single hyperparameter: the rejection percentile threshold. We refer to this procedure as self-normalized resets (SNR) and argue this is a promising approach to mitigating plasticity loss:
@@
    We demonstrate superior performance on four benchmark problems classes studied in (Dohare et al., 2023; Kumar et al., 2023b). Interestingly, there is no single closest competitor to SNR across these problems. Many competing approaches also show significant sensitivity to the choice of hyperparameters; SNR does not. We introduce a new problem to elucidate similar plasticity loss phenomena in the context of language models, and show similar relative merits for SNR.
@@ 1.1 Related Literature
Regularization Methods. L2 regularization has been shown to reduce plasticity loss, but is insufficient in completely alleviating the phenomenon (Dohare et al., 2021; Lyle et al., 2023). While L2 regularization limits weight norm growth during continual learning, it can exacerbate weight rank collapse due to regularization towards the origin. One successful regularization technique is Shrink and Perturb (S&P) (Ash & Adams, 2020), which periodically scales the network’s weights by a shrinkage factor $p$ followed by adding random noise to each weight with scale $\sigma$. Another approach is to perform L2 regularization towards the initial weights referred to as L2 Init (Kumar et al., 2023b). These methods can be viewed as variants of L2 regularization that regularize towards a random initialization and the original initialization, respectively. These methods limit the growth of weight norms while maintaining weight rank and neuron activity by regularizing towards a high-plasticity parameterization.
Architectural and Optimizer Modifications. Architectural modifications such as layer normalization (Ba et al., 2016) and the use of concatenated ReLU activations have been shown to improve plasticity to varying degrees across network architectures and problem settings (Lyle et al., 2023; Kumar et al., 2023b). Additionally, tuning Adam hyperparameters to improve the rate at which second moment estimates are updated has been explored with some success in Lyle et al. (2023).
@@ 2 Algorithm
Plasticity loss describes the phenomenon where, for certain continual learning processes $\Theta_{t}$, such as those corresponding to SGD or Adam, average expected regret increases over time, even for benign choices of $\mu_{t}$. [^1] Specifically, one canonical choice of the sequence of measures $\mu_{t}$ considered in all of the literature on this topic is dividing $T$ into intervals, each of length, say $\Delta$, and having $\mu_{t}$ be constant and equal to $\mu_{i}$ over the $i$th such interval. If $\mu_{i}$ is itself drawn randomly from some distribution of measures and $\Delta$ scales faster than a constant with $T$, we would expect average expected regret to scale like a constant; this is certainly the case if the optimization problem defining $\theta_{t}^{*}$ is convex; in which case that constant is zero. To make these ideas concrete, it is worth considering an example of the above phenomenon reported first by Dohare et al. (2021).
@@
Consider a sequence of ‘tasks’ presented sequentially to SGD, wherein each task consists of 10000 images from the MNIST dataset with the pixels permuted. SGD trains over a single epoch on each task before the subsequent task is presented. Figure 1 measures average accuracy on each task; we see that average accuracy decreases over tasks. The figure also shows a potential correlate of this phenomenon: the number of ‘dead’ or inactive neurons222this notion is formalized in Section 2.1 in the network increases as training proceeds, diminishing the network’s effective capacity.
@@
To motivate our algorithm, SNR, consider applying the network $f(\cdot;\theta_{t}^{*})$ to a hypothetical sequence of training examples drawn i.i.d. from $\mu_{t}$ indexed by $s$. Let $Z^{\mu_{t}}_{s,i}$ indicate the sequence of activations of neuron $i$, and let $A^{\mu_{t}}_{i}$ be a random variable distributed as the random time between any two consecutive activations over this hypothetical sequence of examples. Now turning to the actual sequence of training examples, let $a^{t}_{i}$ count the time since the last firing of neuron $i$ prior to time $t$. Our (idealized) proposal is then exceedingly simple: reset neuron $i$ at time $t$ iff $\mathbb{P}(A_{i}^{\mu_{t}}\geq a^{t}_{i})\leq\eta$ for some suitably small threshold $\eta$. We dub this algorithm Self-Normalized Resets and present it as Algorithm 1. The algorithm requires a single hyper parameter, $\eta$. Of course in practice, the distribution of $A_{i}^{\mu_{t}}$ is unknown to us, and so an implementable version of Algorithm 1 simply approximates this distribution with the histogram of inter-firing times of neuron $i$ prior to time $t$ [^3] As opposed to tracking the histogram itself, we simply track the mean inter-firing time, and assume $A_{i}^{\mu_{t}}$ is geometrically distributed with that mean. This requires tracking just one parameter per neuron. Our mean estimate is itself computed over a fixed length trailing window motivated by the change-point detection approach of Besbes & Zeevi (2011); the length of this window is a hyper-parameter.
@@ 2.1 Motivating SNR and Comparison to Other Reset Schemes
Here we motivate the SNR heuristic and compare it to other proposed reset schemes. Consider the following simple hypothesis test: we observe a discrete time process $Z_{s}\in\{0,1\}$ which under the null hypothesis $H_{0}$ is a Bernoulli process with mean $p>0$. The alternative hypothesis $H_{1}$ is that the mean of the process is identically zero. A hypothesis test must, at some stopping time $\tau$, either reject ($X_{\tau}=1$) or accept ($X_{\tau}=0$) the null; an optimal such test would choose to minimize the sum of type-1 and type-2 errors (the ‘error rate’) and a penalty for delays:
@@
Here the multiplier $\lambda>0$ penalizes the delay in a decision. If $\lambda<p/2$, the optimal test takes a simple form: for some suitable threshold $\bar{T}$, reject the null iff $Z_{s}=0$ for all times $s$ up to $\bar{T}$:
@@
Notice that if $\lambda\propto p$, the percentile threshold above is independent of $p$. Applying this setup to the setting where under the null, we observe the firing of neuron $i$ under i.i.d. training examples from $\mu_{t}$ and a neuron is considered ‘dead’ or inactive if the alternate hypothesis is true, imagine that $p=\mathbb{P}(Z^{\mu_{t}}_{s,i}=1)$. Further, we assume $\lambda=\alpha p$ ($\alpha<1/2$); a reasonable assumption which models a larger penalty for late detection of neurons that are highly active. It is then optimal to declare neuron $i$ ‘inactive’ if the length of time it has not fired exceed the $1-\alpha(1-\alpha)^{-1}$ percentile of the distribution of $A_{i}^{\mu_{t}}$. This is the underlying motivation for the SNR heuristic.
@@
This reveals an interesting comparison with SNR. The schemes above will re-initialize a neuron after inactivity over a period of time that is uniform across all neurons. On the other hand, SNR will reset a neuron after it is inactive for a period that corresponds to a fixed percentile of the inter-firing time distribution of that neuron. Whereas this percentile is fixed across neurons, the corresponding period of inactivity after which a neuron is reset will vary across neurons: shorter for neurons that tend to fire frequently, and longer for neurons that fire less frequently.
We can make this comparison precise in the context of the hypothesis testing setup above: specifically, consider two neurons with null firing rates $p_{1}$ and $p_{2}$ respectively ($p_{1}<p_{2}$), and delay multipliers, $\lambda$, of $\alpha p_{1}$ and $\alpha p_{2}$ respectively. By Proposition 2.1, under SNR, the first is reset if it is inactive for time at least $\log(\alpha(1-\alpha)^{-1})/\log(1-p_{1})$ and the second if it is inactive for time at least $\log(\alpha(1-\alpha)^{-1})/\log(1-p_{2})$. In contrast, for a fixed threshold scheme such as Sokar et al. (2023), either neuron would be reset after being inactive for some fixed threshold, say $r^{*}$. Assume $r^{*}$ is set to minimize the sum of the error rates of the two neurons while keeping the total delay identical to that for SNR. The proposition below compares the error rate between the two schemes:
Proposition 2.2.
The ratio of total error rate with a fixed threshold $r^{*}$ to that under SNR scales like
$\Omega\left(\exp\left(\log(\alpha(1-\alpha)^{-1})\left(-\frac{1}{2}+\frac{1}{2}\frac{\log(1-p_{1})}{\log(1-p_{2})}\right)\right)\right)$
Now recall that $\alpha<1/2$ and $p_{1}<p_{2}$. The result above then shows that: (a) the error rate under an (optimal) fixed threshold can grow arbitrarily larger than the error rate under SNR as the penalty for delay shrinks to zero and (b) the rate at which this gap grows itself scales with the difference in the nominal firing rates of the neurons under consideration. This provides insight into the relative merits of using a scheme like SNR in lieu of existing reset proposals: resets under SNR detect changes in the firing rate of a neuron faster and more accurately; this matters particularly in situations where there is wide disparity in the nominal firing rates of neurons across the network.
@@ 3.1 Experimental Setup
Permuted MNIST (PM) (Goodfellow et al., 2013; Dohare et al., 2021; Kumar et al., 2023b): A subset of 10000 image-label pairs from the MNIST dataset are sampled for an experiment. A task consists of a random permutation applied to each of the 10000 images. The network is presented with 2400 tasks appearing in consecutive order. Each task consists of a single epoch and the network receives data in batches of size 16.
Random Label MNIST (RM) (Kumar et al., 2023b; Lyle et al., 2023): A subset of 1200 images from the MNIST dataset are sampled for an experiment. An experiment consists of 100 tasks, where each tasks is a random assignment of labels, consisting of 10 classes, to the 1200 images. A network is trained for 400 epochs on each task with a batch size 16.
Random Label CIFAR (RC) (Kumar et al., 2023b; Lyle et al., 2023): A subset of 128 images from the CIFAR-10 dataset are sampled for an experiment. An experiment consists of 50 tasks, where each tasks is a random assignment of labels, consisting of 10 classes, to the 128 images. An agent is trained for 400 epochs on each task with a batch size 16.
Continual Imagenet (CI) (Dohare et al., 2023; Kumar et al., 2023b): An experiment consists of all 1000 classes of images from the ImageNet-32 dataset (Chrabaszcz et al., 2017) containing 600 images from each class. Each task is a binary classification problem between two of the 1000 classes, selected at random. The experiment consists of 500 tasks and each class occurs in exactly one task. Each task consists of 1200 images, 600 from each class, and the network is trained for 10 epochs with a batch size of 100.
@@
This experimental setup, for all but Permuted Shakespeare, follows that of (Kumar et al., 2023b), with the exceptions of Permuted MNIST which has its task count increased from 500 to 2400, Random Label MNIST which has its task count increased from 50 to 100, and Random Label CIFAR which has its dataset reduced from 1200 to 128 images. Lyle et al. (2023) consider variants of the Random Label MNIST and CIFAR problems by framing them as MDP environments for DQN agents. During training, the DQN agents are periodically paused to assess their plasticity by training them on separate, randomly generated regression tasks using the same image datasets.
@@ 3.1.1 Algorithms and Architetcures
Our baseline in all problems consist simply of using SGD or Adam as the optimizer with no further intervention. We then consider several interventions to mitigate plasticity loss. First, we consider algorithms that employ an explicit reset of neurons: these include SNR, Continual Backprop (CBP) (Dohare et al., 2021), and ReDO (Sokar et al., 2023). Among algorithms that attempt to use regularization, we consider vanilla L2 regularization, L2 Init (Kumar et al., 2023b), and Shrink and Perturb (Ash & Adams, 2020). Finally, as a potential architectural modification we consider the use of Layer Normalization (Ba et al., 2016).
@@
MLP: For Permuted MNIST and Random Label MNIST we use an MLP identical to that in Kumar et al. (2023b) which in turn is a slight modification to that in Dohare et al. (2023).
CNN: For Random Label CIFAR and Continual ImageNet we use a CNN architectures identical to that in Kumar et al. (2023b) which in turn is a slight modification to that in Dohare et al. (2023).
@@ 3.2 Results and Discussion
Table 1: Average accuracy on the last 10% of tasks on the benchmark continual learning problems over 5 seeds. Standard deviations are provided in the extended Table 6.
@@
L2 $0.2762\hskip 5.0pt(0.0309)$ $0.1560\hskip 5.0pt(0.0236)$ $0.3101\hskip 5.0pt(0.0425)$
SNR+L2 $0.2177\hskip 5.0pt(0.0196)$ $0.1370\hskip 5.0pt(0.0169)$ $0.2551\hskip 5.0pt(0.0454)$
No Intv. $2.7397\hskip 5.0pt(0.0140)$ $1.8164\hskip 5.0pt(0.0295)$ $3.0147\hskip 5.0pt(0.0250)$
L2 Init $1.5052\hskip 5.0pt(0.0437)$ $1.2931\hskip 5.0pt(0.0486)$ $1.5262\hskip 5.0pt(0.0420)$
SNR $2.6872\hskip 5.0pt(0.0222)$ $1.7338\hskip 5.0pt(0.0295)$ $3.0242\hskip 5.0pt(0.0408)$
CBP 2.4922    (0.0171) 1.3732    (0.0234) 2.9410    (0.0188)
L2* 0.1506    (0.0279) 0.1874    (0.0348) 0.1092    (0.0429)
SNR+L2* 0.1402    (0.0246) 0.1549    (0.0107)    0.0909    (0.0177)
ReDO 2.3258    (0.0356) 1.3689    (0.0686) 2.8119    (0.0373)
Table 2: Average loss measured on the final epoch of each task with standard deviations over 9 seeds on the Permuted Shakespeare problem. Note, L2* denotes L2 regularization applied only to the attention weights.
@@
While L2 regularization addresses weight blowup, neuron death remains present; see Figure 3. The average loss with L2 increases from 0.1560, over the first 50 tasks, to 0.3101, over the final 50 tasks. This prompts us to consider using SNR in addition to L2 regularization. This largely eliminates neuron death (see the right panel of Figure 3), while stabilizing weight norms and maintaining entropy of self-attention probabilities, providing the lowest loss (0.2551) over the final 50 tasks. As an ablation, we evaluate performance when applying L2 regularization only to the attention weights, which we denote by L2* and SNR+L2* in Table 2.
Figure 3: Number of inactive neurons for 5 consecutive training steps on the PS experiment.